OpenStudy (anonymous):

The Weight of heavy water containing 5*10^5 D atoms is?

OpenStudy (zepp):

D atoms?

OpenStudy (anonymous):

\[\frac{5e5}{6.022e23}*2\]

OpenStudy (anonymous):

did nt understand..

OpenStudy (anonymous):

Find no of moles and multiply by gram molar mass

OpenStudy (unklerhaukus):

\[\frac{5\times10^5atomsD}{6.02\times10^{23}atoms /mol}\times \frac{2g/mol}{D}=1.66\times10^{-18}g\]

OpenStudy (anonymous):

the answer should be 8.3* 10^-18g

OpenStudy (unklerhaukus):

that is for regular hydrogen @shameer1

OpenStudy (anonymous):

did nt understand

OpenStudy (anonymous):

@.Sam. @myininaya @blues

sam (.sam.):

If its deuterium then Unkle is right

OpenStudy (unklerhaukus):

actually now i think i have miss interpreted the question the question is asking for mass of heavy water, not a mass of Deuterium, heavy water has a mixture of normal hydrogen and deuterium in the water molecules

OpenStudy (anonymous):

can anyone answer it now..

OpenStudy (blues):

Call me an idiot if I am wrong but I believe the molar mass of heavy water is about 20 g / mol, not 2 as above...

OpenStudy (unklerhaukus):

yeah i was getting confused with the hydrogen and the water

OpenStudy (anonymous):

molar mass of h20 = 18 is it,....

sam (.sam.):

H-2 is 2 g/mol, checked, http://www.ptable.com/